OpenStudy (anonymous):

Solve the equation: (x-3)(x+2)=0

OpenStudy (anonymous):

x^2-x-6

OpenStudy (anonymous):

it needs to be solved from there as well

OpenStudy (anonymous):

how so?

myininaya (myininaya):

(x-3)(x+2)=0 x-3=0,x+2=0 x=3,x=-2

OpenStudy (anonymous):

ohhh you want x values? x-3=0 x=3 x+2=0 x=-2

OpenStudy (anonymous):

yes, thank you